OpenStudy (anonymous):

A rectangular pizza is cut into 6 pieces along the length and 4 pieces along the width. If you place 3 tomato slices on each piece of pizza, how many tomato slices will you use? A. 8 B. 72 C. 60 D.24 please help !

OpenStudy (anonymous):

6*4=24 3*24=72

OpenStudy (anonymous):

Thank you ! (:

OpenStudy (anonymous):

i'm so hungry . i want a piece of this delicious pizza

OpenStudy (apoorvk):

(no. of pieces) * (no. of tomato slices per piece) = 6 * 3 = ?? P.S.- olives would have been better :P

OpenStudy (anonymous):

6x3 = 18

OpenStudy (apoorvk):

lol, and also multiply by '4', for the no. of pieces width wise - i missed that bit.

OpenStudy (anonymous):

So i multiply 6x3x4 ?

OpenStudy (anonymous):

@apoorvk : i agree ; olive is better . i want it now :D

OpenStudy (anonymous):

@apoorvk ?

OpenStudy (apoorvk):

yeah multiply the three!!

OpenStudy (anonymous):

72 ...

OpenStudy (apoorvk):

oops, what a bug.. yeah, 72 is right! that should call for some pizza, @Roya and @xoxosweetx3 http://cdn.phamfatale.com/album/olive-and-tomato-cheese-pizza.jpg

OpenStudy (anonymous):

@apoorvk ; kind of u dear .

OpenStudy (apoorvk):

@Roya , you see, I am sooooooo kind!! :D lol

OpenStudy (anonymous):

@apoorvk : of course u r dear .

OpenStudy (apoorvk):

haha, you too!!